# STRIDE: Single-video based Temporally Continuous Occlusion Robust 3D Pose Estimation

### v2

@@ 3.2 Learning a Motion Prior
To build a strong prior for human motion dynamics, we draw inspiration from the success of large language models like BERT [9] that leverage large-scale self-supervised pre-training. Here, we extend this paradigm to 3D human pose estimation. Specifically, given a dataset of 3D pose sequences, we synthetically mask these sequences to simulate occlusions and other errors. Similar to [7, 52], the prior $\mathcal{M}$ is trained to denoise these noisy inputs to reconstruct a sequence of temporally coherent 3D poses.
During pre-training, we apply both joint-level and frame-level masking to a 3D pose sequence $\mathbf{X}$ to obtain a corrupted sequence $\texttt{mask}(\mathbf{X})$ which mimics realistic scenarios of imperfect predictions and occlusions. The prior $\mathcal{M}$ is trained to reconstruct the complete 3D motion sequence $\mathbf{\bar{X}}$ from this corrupted input $\mathbf{{X}}$ by minimizing losses on 3D joint positions $\mathcal{L}_{\text{3D}}$ between the reconstruction and the ground-truth pose. Additionally, we incorporate a velocity loss $\mathcal{L}_{\text{O}}$ following [34, 45].
$\mathcal{L}_{\text{3D}}=\sum\limits_{t=1}^{T}\sum\limits_{j=1}^{J}\parallel\mathbf{\bar{X}}_{t,j}-\mathbf{X}_{t,j}\parallel_{2}\qquad\mathcal{L}_{\text{O}}=\sum\limits_{t=2}^{T}\sum\limits_{j=1}^{J}\parallel\mathbf{\bar{O}}_{t,j}-\mathbf{O}_{t,j}\parallel_{2}$
where $\mathbf{\bar{O}}_{t}=\mathbf{\bar{{X}}}_{t}-\mathbf{\bar{X}}_{t-1}$, $\mathbf{O}_{t}=\mathbf{X}_{t}-\mathbf{X}_{t-1}$.
@@ 3.3 Test-Time Alignment
Although the prior refines pose, some inconsistencies like domain shift and novel human motion may be present in the videos. Hence, we propose additional test-time training of $\mathcal{M}$ using geometric and physics-based constraints to adapt to such situations. Similar to internal learning approaches like Deep Video Prior [23], our proposed self-supervision strategy fine-tunes the motion prior to the specifics of each test video for enhanced outputs. Specifically, we utilize four different losses that regularize (1) the velocity of joints, (2) scale variations in predictions (3) the size of limbs, and (4) the smoothness of poses (5) in missing frames. Crucially, only $\mathcal{M}$ is updated during test-time training while $\mathcal{P}$ remains fixed to preserve the pose estimation capabilities of off-the-shelf models.
@@
$\mathcal{L}_{\text{NMPJP}}=\mathcal{L}_{\text{MPJP}}(s\mathbf{\hat{X}},\mathbf{\tilde{X}});\;\text{where }s=\frac{\sum_{t=1}^{T}\sum_{j=1}^{J}\sum_{d=1}^{3}\left\|\mathbf{\tilde{X}}_{t,j,d}\cdot\mathbf{\hat{X}}_{t,j,d}\right\|_{2}}{\sum_{t=1}^{T}\sum_{j=1}^{J}\sum_{d=1}^{3}\left\|\mathbf{\hat{X}}_{t,j,d}\right\|_{2}^{2}}$ (3)
In Equation 3, $s$ represents the scale. The combination of both $\mathcal{L}_{\text{NMPJP}}$ and $\mathcal{L}_{\text{MPJP}}$ losses allows the model to simultaneously optimize for accurate joint positions ($\mathcal{L}_{\text{MPJP}}$) and address scale variations ($\mathcal{L}_{\text{NMPJP}}$). The incorporation of $\mathcal{L}_{\text{NMPJP}}$ allows the model to learn to handle scenarios where the pose scale may differ between training and testing data. Velocity loss: We optimize velocity loss similar to Eq. 4, but instead of ground truth, we use pseudo-labels,
@@
Overall Loss. In summary, by combining all the above-mentioned losses into one final loss function as shown in Eq. 5, $\mathcal{M}$ is trained to produce accurate joint positions, maintain anatomical consistency, and handle scale variations,
$\mathcal{L}_{total}=\lambda_{1}\mathcal{L}_{mpjp}+\lambda_{2}\mathcal{L}_{vel}+\lambda_{3}\mathcal{L}_{lim}+\lambda_{4}\mathcal{L}_{nmpjp}$ (5)
@@ 4 Experiments and Results
We conduct evaluations on three datasets with varying levels of occlusion: Human3.6M, representing scenarios without occlusion; OCMotion, moderate occlusion; and Occluded Human3.6M, representing heavy occlusion. The metrics assessed include Procrustes-aligned mean per joint position error (PA-MPJPE), mean per joint position error (MPJPE), and acceleration error (Accel), measured as the disparity in acceleration between ground-truth and predicted 3D joints. We report the metrics in (mm). We use BEDLAM-CLIFF [3] as the off-the-shelf pose estimation method. We compare the error rates of STRIDE  and the baseline methods in Table 1,2 and 3. The best results are in bold and green arrows indicate the percentage improvement over the previous state-of-the-art method.
@@ 4.1 Datasets
Human3.6M [15]. An indoor-scene dataset, Human3.6M is a pivotal benchmark for 3D human pose estimation from 2D images. Captured with a 4-camera setup, it includes 11 subjects, each with 15 different actions, annotated in the 17 keypoints format. Following [3], we retain every 1 in 5 frames in the test split comprising the S9 and S11 sequence. We perform experiments on the original publically available Human3.6M dataset to show that our method achieves comparable performance with other state-of-the-art methods.
OCMotion [14]. OCMotion is a video dataset that extends the 3DOH50K image dataset [46], incorporating natural occlusions. The dataset comprises 300K images captured at 10 FPS, featuring 43 sequences observed from 6 viewpoints. Its annotations for 3D motion include SMPL, 2D poses, and camera parameters. The sequences {0013, 0015, 0017, 0019} are designated for testing. Our method does not require supervised training, so we have only used the test split when performing all experiments.
Occluded Human3.6M. We curate the Occluded Human3.6M dataset to evaluate our method, specifically designed for assessing human pose estimation under significant occlusion, unlike existing datasets such as Human3.6M, MPI-INF-3DHP[28], and 3DPW[25]. To accomplish this, we use random erase occlusions on Human3.6M videos, completely covering a person up to 100%. These occlusions persist spatially and temporally for 1.6 seconds within 3.2 seconds of the video. Note: Additional datasets, baselines and implementation details can be found in the suppl. material.
BRIAR [8]. BRIAR is a large-scale biometric dataset featuring videos of human subjects captured in extremely challenging conditions. These videos are recorded at varying distances $i.e$ close range, 100m, 200m, 400m, 500m, and unmanned aerial vehicles (UAV), with each video lasting around 90 seconds. Most of the pose estimation methods fail on this dataset due to the extreme amount of domain shifts. Additionally, BRIAR lacks ground truth data for poses, which means evaluations of pose estimation methods on this dataset can only be qualitative, relying on visual assessments rather than quantitative metrics.
@@ 4.2 Quantitative Results
Our method is most effective under heavy occlusions. We significantly outperform other state-of-the-art methods on the Occluded Human3.6M dataset as shown in Table 1. Notably, STRIDE  performs significantly better than BEDLAM despite using pseudo-labels from BEDLAM. BEDLAM fails to produce poses under heavy occlusion; hence, the evaluation results drop significantly. However, since STRIDE  incorporates temporal information to address these gaps in the video, we predict reasonable poses even in case of heavy occlusions and improve the result of BEDLAM by a significant margin. It is important to note that by using STRIDE  we do not only outperform BEDLAM, but we also outperform all the other existing video- and image-based state-of-the-art methods. This is mainly because existing methods do not incorporate human motion prior and hence results in temporally implausible poses.
Method PA-MPJPE Accel Avg Image OOH [46] 55.0 48.6 51.8 PARE [18] 52.0 43.6 47.8 BEDLAM [3] 47.1 49.0 48.0 Video PoseFormerV2  [47] 126.3 28.5 77.4 GLAMR  [43] 89.9 51.3 70.6 CycleAdapt  [30] 74.6 57.5 66.0 ROMP [40] 48.1 57.2 52.6 STRIDE (ours) 46.2 (2$\downarrow$) 47.8 47.0 (2$\downarrow$) Table 2: 3D pose estimation results on OCMotion [14]. STRIDE  outperforms other image and video-based pose estimation methods. While PoseFormerV2 has the lowest accel., it also exhibits the highest PA-MPJPE error. This is due to oversmoothing and inaccurate interpolation between poses which compromises the pose estimation accuracy.
Since Occluded Human3.6M contains artificial occlusions, we also evaluated on the OCMotion dataset, which contains real-world, natural occlusions. Table 2 shows that our approach STRIDE attains state-of-the-art results on the OCMotion dataset [14]. Since we obtained good pseudo-labels from BEDLAM under partial occlusions, we observe the proximity of our results to BEDLAM. It is important to highlight that methods such as [40, 18] are supervised and trained on the training split of OCMotion. In contrast, our approach does not assume access to any labeled training dataset.
@@
Table 3: 3D pose estimation results on Human3.6M. Our evaluation demonstrates that our results are comparable to the BEDLAM-CLIFF baseline. This is due to the occlusion-free nature of the Human3.6M, which yields already refined and consistent poses with limited room for improvement
Our method demonstrates minor improvement over BEDLAM-CLIFF on the original Human3.6M dataset, as evidenced in Table 3. The marginal enhancement is primarily due to the nature of the Human3.6M dataset, which lacks occlusions, thereby limiting the potential for improvement beyond the baseline. A thorough analysis of our findings, including the observed enhancement in temporal smoothness, is provided in the suppl. material.
Inference speed: Table  4 compares the inference times of various 3D pose
@@
estimation methods on a 243-frame OCMotion video using an RTX 3090 GPU. HuMor and GLAMR are notably slower, exceeding 10 minutes due to their intensive pose optimization phase. In contrast, PoseFormerV2 and CycleAdapt show efficiency improvements with inference times of 129 and 126 seconds, respectively. STRIDE  outperforms these, achieving a significant reduction to 68 seconds, making it 46% faster and highlighting its suitability for real-time applications without sacrificing accuracy.
@@ 4.3 Qualitative Results
To provide a comprehensive analysis and comparison of STRIDE  against other methods, we have compiled and shared several qualitative video results in the suppl. material. Our evaluation juxtaposes STRIDE  against leading state-of-the-art techniques like CycleAdapt [30]. Key insights from our comparison include:
Occluded Human3.6M: Traditional approaches often fall short in accurately predicting missing 3D poses, struggling with high levels of occlusion. In contrast, our method utilizes the dynamics of human motion to precisely infill missing poses, leading to a 57% error improvement in performance compared to the former state-of-the-art method. These improvements can be visualized in Fig. 4.
Figure 4: 3D pose estimation results on Occluded Human3.6M. CycleAdapt (second row) fails to generalize in cases when there is complete occlusion. STRIDE  (third row) produces temporally coherent pose infilling due to test time training. Note that the translucent red color represents the ground truth poses.
@@
Refer to the suppl. material for additional qualitative 3D pose estimation, details on mesh generation in videos and mesh recovery results.
@@ 4.4 Ablation Study
An ablation study conducted in Table 5 provides quantitative insights into the significance of each component in STRIDE. Starting from a baseline with substantial errors, the introduction of a motion prior alone drastically improves performance, underscoring its effectiveness in driving the model toward realistic human pose dynamics. The addition of $L_{mpjp}$ enhances spatial accuracy, further lowering MPJPE to 82.1 and PA-MPJPE to 60.4. The improvement with $L_{vel}$ suggests its role in smoothing motion. The best results are observed when $L_{nmpjp}$ is also included, indicating its critical function in accounting for scale variations.
In conclusion, the ablation study reveals that each component contributes to improving the accuracy and temporal consistency of the pose estimations, with the full combination of components yielding the state-of-the-art results. This indicates that while the motion prior sets a strong foundation for plausible poses, the various loss functions refine and stabilize the pose predictions to align closely with natural human movement dynamics and unseen poses.
Prior $\mathcal{L}_{mpjp}$ $\mathcal{L}_{vel}$ $\mathcal{L}_{lim}$ $\mathcal{L}_{nmpjp}$ MPJPE PA-MPJPE
@@
Table 5: Ablation study This table demonstrates how the inclusion of a pre-trained motion prior and various loses collectively contributes to the model’s accuracy on Occluded Human3.6M dataset.
In the suppl. material, we demonstrate how varying off-the-shelf pose estimation methods within the backbone of STRIDE  affects its performance. We find that using any off-the-shelf pose estimation method yields similar improvements, thereby making STRIDE  agnostic to any specific 3D pose estimation method.
5 Conclusion and Future Works
In conclusion, while existing 3D human pose estimation methods excel in various scenarios, they struggle with handling significant occlusions. In this work, we introduce STRIDE, an unsupervised approach which utilizes large-scale pre-training, self-supervised learning, and temporal context to enhance 3D pose estimation for a single video containing occlusions during the test time. STRIDE achieves state-of-the-art results on datasets that contain significant human body occlusions such as Occluded Human3.6M and OCMotion thus demonstrating improved occlusion robustness. Currently, a limitation of STRIDE  is that it can only extract temporally continuous 3D poses when there are no human-to-human occlusions. Future work will focus on adapting STRIDE for multi-person occlusion scenarios. Future work can also involve the using temporally continuous pose estimates to enhance downstream tasks such as action recognition, mesh recovery, and gait recognition.
6 Acknowledgements
This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via [2022-21102100007]. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.

### v4

@@ 3.2 Learning a Motion Prior
To build a strong prior for human motion dynamics, we draw inspiration from the success of large language models like BERT [9] that leverage large-scale self-supervised pre-training. Here, we extend this paradigm to 3D human pose estimation. Specifically, given a dataset of 3D pose sequences, we synthetically mask these sequences to simulate occlusions and other errors. Similar to [7, 51], the prior $\mathcal{M}$ is trained to denoise these noisy inputs to reconstruct a sequence of temporally coherent 3D poses. We selected BERT-style training for STRIDE’s motion prior because it is highly effective at capturing bidirectional dependencies in human motion data and excels in scenarios where pose information may be incomplete or corrupted. BERT-style training leverages a bidirectional approach, allowing the model to consider both preceding and succeeding context simultaneously. This is particularly advantageous for understanding human motion, as it mirrors real-world situations where poses might be partially missing or noisy.
During pre-training, we apply both joint-level and frame-level masking to a 3D pose sequence $\mathbf{X}$ to obtain a corrupted sequence $\texttt{mask}(\mathbf{X})$ which mimics realistic scenarios of imperfect predictions and occlusions. The prior $\mathcal{M}$ is trained to reconstruct the complete 3D motion sequence $\mathbf{\bar{X}}$ from this corrupted input $\mathbf{{X}}$ by minimizing losses on 3D joint positions $\mathcal{L}_{\text{3D}}$ between the reconstruction and the ground-truth pose. Additionally, we incorporate a velocity loss following [30, 44].
@@ 3.3 Test-Time Alignment
Although the prior refines pose, some inconsistencies like domain shift and novel human motion may be present in the videos. Hence, we propose additional test-time training of $\mathcal{M}$ using geometric and physics-based constraints to adapt to such situations. Similar to internal learning approaches like Deep Video Prior [23], our proposed self-supervision strategy fine-tunes the motion prior to the specifics of each test video for enhanced outputs. We use four loss regularizers targeting different aspects of human motion: (1) Limb Loss, (2) Mean Per Joint Position (MPJP) Loss, (3) Normalized MPJP (N-MPJP) Loss, and (4) Velocity Loss. Crucially, only $\mathcal{M}$ is updated during test-time training while $\mathcal{P}$ remains fixed to preserve the pose estimation capabilities of off-the-shelf models.
@@
$\mathcal{L}_{\text{NMPJP}}=\mathcal{L}_{\text{MPJP}}(s\mathbf{\hat{X}},\mathbf{\tilde{X}})$ (3)
where $\text{where }s=\frac{\sum_{t=1}^{T}\sum_{j=1}^{J}\sum_{d=1}^{3}\left\|\mathbf{\tilde{X}}_{t,j,d}\cdot\mathbf{\hat{X}}_{t,j,d}\right\|_{2}}{\sum_{t=1}^{T}\sum_{j=1}^{J}\sum_{d=1}^{3}\left\|\mathbf{\hat{X}}_{t,j,d}\right\|_{2}^{2}}$ (4)
In Equation 3, $s$ represents the scale. The combination of both $\mathcal{L}_{\text{NMPJP}}$ and $\mathcal{L}_{\text{MPJP}}$ losses allows the model to simultaneously optimize for accurate joint positions ($\mathcal{L}_{\text{MPJP}}$) and address scale variations ($\mathcal{L}_{\text{NMPJP}}$). The incorporation of $\mathcal{L}_{\text{NMPJP}}$ allows the model to learn to handle scenarios where the pose scale may differ between training and testing data. Velocity Loss: We optimize velocity loss similar to Equation 5, but instead of ground truth, we use pseudo-labels,
@@
Overall Loss. In summary, by combining all the above-mentioned losses into one final loss function as shown in Equation 6, $\mathcal{M}$ is trained to produce accurate joint positions, maintain anatomical consistency, and handle scale variations,
$\mathcal{L}_{total}=\lambda_{1}\mathcal{L}_{\text{lim}}+\lambda_{2}\mathcal{L}_{\text{MPJP}}+\lambda_{3}\mathcal{L}_{\text{NMPJP}}+\lambda_{4}\mathcal{L}_{\text{vel}}$ (6)
@@ 4 Experiments and Results
We conduct evaluations on three datasets with varying levels of occlusion: Human3.6M, representing scenarios without occlusion; OCMotion, moderate occlusion; and Occluded Human3.6M, representing heavy occlusion. The metrics assessed include Procrustes-aligned mean per joint position error (PA-MPJPE), mean per joint position error (MPJPE), and acceleration error (Accel), measured as the disparity in acceleration between ground-truth and predicted 3D joints. We report the metrics in (mm). We use BEDLAM-CLIFF [3] as the off-the-shelf pose estimation method. We compare the error rates of STRIDE  and the baseline methods in Tables 1, 2 and 3. The best results are in bold and arrows indicate the percentage improvement over the best existing algorithm. Qualitative video results can be found on our shared GitHub repository.
@@ 4.1 Datasets
Human3.6M [16]: This indoor-scene dataset is crucial for 3D human pose estimation from 2D images. We use every 1 in 5 frames in the test split and achieve comparable performance to state-of-the-art methods.
OCMotion [15]: This video dataset extends the 3DOH50K image dataset with natural occlusions, comprising 300K images at 10 FPS. We use only the test split since our method does not require supervised training.
Occluded Human3.6M: We prepare a new dataset by modifying the Human3.6M [16]. It is curated to evaluate pose estimation under significant occlusion. This dataset uses random erase occlusions covering up to 100% of a person for 1.6 seconds within 3.2-second videos.
BRIAR [8]: Features videos of human subjects in extremely challenging conditions, recorded at varying distances and from UAVs. Additional details on datasets and implementation specifics, are provided in Section 7 of supplementary material.
@@ 4.2 Quantitative Results
Our method is most effective under heavy occlusions. We significantly outperform other state-of-the-art methods on the Occluded Human3.6M dataset as shown in Table 1. Notably, STRIDE  performs significantly better than BEDLAM despite using pseudo-labels from BEDLAM. BEDLAM fails to produce poses under heavy occlusion; hence, the evaluation results drop significantly. However, since STRIDE  incorporates temporal information to address these gaps in the video, we predict reasonable poses even in case of heavy occlusions and improve the result of BEDLAM by a significant margin. It is important to note that by using STRIDE  we do not only outperform BEDLAM, but we also outperform all the other existing video- and image-based state-of-the-art methods. This is mainly because existing methods do not incorporate human motion prior and hence result in temporally implausible poses.
Method PA-MPJPE Accel Avg
Image OOH [45] 55.0 48.6 51.8
PARE [19] 52.0 43.6 47.8
BEDLAM [3] 47.1 49.0 48.0
Video PoseFormerV2 [47] 126.3 28.5 77.4
GLAMR [41] 89.9 51.3 70.6
CycleAdapt [26] 74.6 57.5 66.0
ROMP [37] 48.1 57.2 52.6
STRIDE (ours) 46.2 (2$\downarrow$) 47.8 47.0 (2$\downarrow$)
Table 2: 3D pose estimation results on OCMotion [15]. STRIDE  outperforms other image and video-based pose estimation methods. While PoseFormerV2 has the lowest accel., it also exhibits the highest PA-MPJPE error. This is due to oversmoothing and inaccurate interpolation between poses which compromises the pose estimation accuracy.
Since Occluded Human3.6M contains artificial occlusions, we also evaluated on the OCMotion dataset, which contains real-world, natural occlusions. Table 2 shows that our approach STRIDE attains state-of-the-art results on the OCMotion dataset [15]. Since we obtained good pseudo-labels from BEDLAM under partial occlusions, we observe the proximity of our results to BEDLAM. It is important to highlight that methods such as [37, 19] are supervised and trained on the training split of OCMotion. In contrast to these algorithms, our proposed approach (STRIDE) does not assume access to any labeled training dataset.
@@
Table 3: 3D pose estimation results on Human3.6M. Our evaluation demonstrates that our results are comparable to the BEDLAM-CLIFF baseline. This is due to the occlusion-free nature of the Human3.6M, which yields already refined and consistent poses with limited room for improvement.
Our method demonstrates minor improvement over BEDLAM-CLIFF [3] on the original Human3.6M dataset, as evidenced in Table 3. The marginal enhancement is primarily due to the nature of the Human3.6M dataset, which lacks occlusions, thereby limiting the potential for improvement beyond the baseline. A thorough analysis of our findings, including the observed enhancement in temporal smoothness, is provided in the Section 3 of supplementary.
Inference speed: Table 4 compares the inference times of various 3D pose estimation methods on a 243-frame OCMotion video using an RTX 3090 GPU. HuMor and GLAMR are notably slower, exceeding 10 minutes due to their intensive pose optimization phase. In contrast, PoseFormerV2 and CycleAdapt show efficiency improvements with inference times of 129 and 126 seconds, respectively. STRIDE  outperforms these, achieving a significant reduction to 68 seconds, making it 46% faster and highlighting its suitability for real-time applications without sacrificing accuracy. We have ensured that the testing is consistent across all algorithms, and results are reported including the TTT phase with 30 epochs. Our method takes less time because we train only on a single video at test time, which contains significantly fewer frames compared to the entire dataset.
@@ 4.3 Qualitative Results
To provide a comprehensive analysis and comparison of STRIDE  against other methods, we have compiled and shared several qualitative video results in the supplementary material. Our evaluation juxtaposes STRIDE  against leading state-of-the-art techniques like CycleAdapt [26]. Key insights from our comparison include:
Occluded Human3.6M: In this proposed dataset, traditional approaches often fall short in accurately predicting missing 3D poses, struggling with high levels of occlusion. In contrast, our method utilizes the dynamics of human motion to precisely infill missing poses, leading to a 57% error improvement compared to the former methods.
@@
Please refer to the Section 4 supplementary material for additional qualitative 3D pose estimation results and Section 6 of supplementary material for details on mesh generation in videos and mesh recovery results.
@@ 4.4 Ablation Study
$\mathcal{M}$ $\mathcal{L}_{mpjp}$ $\mathcal{L}_{vel}$ $\mathcal{L}_{lim}$ $\mathcal{L}_{nmpjp}$ MPJPE PA-MPJPE
@@
Table 5: Ablation study. This table illustrates how integrating a pre-trained motion prior and various losses collectively contribute to STRIDE’s final accuracy on the Occluded Human3.6M dataset.
An ablation study conducted in Table 5 provides insights into the significance of each component in STRIDE. Starting from a baseline with substantial errors, introducing a motion prior alone drastically improves performance, underscoring its effectiveness in driving the model toward realistic human pose dynamics. Adding $L_{mpjp}$ enhances spatial accuracy, further lowering MPJPE to 82.1 and PA-MPJPE to 60.4. Improvement with $L_{vel}$ suggests its role in smoothing motion. The best results are observed when $L_{nmpjp}$ is also included, indicating its critical function in accounting for scale variations. Thus, the ablation study reveals that each component contributes to improving the accuracy and temporal consistency of the pose estimations, with the full combination of components yielding state-of-the-art results. This shows that while the motion prior sets a strong foundation for plausible poses, the various loss functions refine and stabilize the pose predictions to align closely with natural human movement dynamics and unseen poses. In Section 5 of supplementary material, we show how varying off-the-shelf pose estimation methods within the backbone of STRIDE affects its performance. We find that using any off-the-shelf pose estimation method yields similar improvements, thereby making STRIDE agnostic to any specific 3D pose estimation method.
5 Conclusion
We introduce STRIDE, a novel algorithm for self-supervised test-time training aimed at improving 3D human pose estimation in individual video frames. STRIDE utilizes extensive self-supervised pre-training to develop a robust model of human motion priors. It integrates self-supervised optimization with temporal regularization, achieving state-of-the-art performance in terms of both pose accuracy and computational efficiency across diverse challenging datasets, even those with significant occlusions. A limitation of STRIDE is its ability to extract temporally continuous 3D poses only in scenarios where there are no human-to-human occlusions. Future efforts can concentrate on adapting STRIDE  for situations involving multi-person occlusions. Handling such situations requires modelling of complex human-to-human interaction alongside external identification and tracking framework. In conclusion, STRIDE sets a new benchmark for 3D human pose estimation in occluded environments and introduces promising directions for enhancing related downstream applications.
Acknowledgment
The work was partially supported by NSF grants 2326309 and 2312395, DURIP grant N000141812252, USDA grant 2021-67022-33453, and the Office of the Director of National Intelligence (ODNI), specifically through the Intelligence Advanced Research Projects Activity (IARPA), under contract number [2022- 21102100007]. The views and conclusions in this research reflect those of the authors and should not be construed as officially representing the policies, whether explicitly or implicitly, of ODNI, IARPA, or the U.S. Government. Nevertheless, the U.S. Government retains the authorization to reproduce and distribute reprints for official government purposes, regardless of any copyright notices included.